OpenStudy (anonymous):

Solve for x. 12 – 3x – x = 8 A. –10 B. –5 C. 1 D. 2

OpenStudy (gorv):

12-3x-x=8 12-4x=8 12-8=4x 4=4x x=1

OpenStudy (anonymous):

Identify the solution set of the given inequality using the given replacement set. x > –4; {–8, –4.3, –4 –3.8, 1.2, 3} A. {–8, –4.3, –4} B. {–8, –4.3} C. {–4, –3.8, 1.2, 3} D. {–3.8, 1.2, 3}